# Deep Learning for Camera Calibration and Beyond: A Survey

### v2

@@ 3.4.2 Future Effort
(1) Explore more model priors. Most learning-based calibration methods study the parametric camera models but their generalization abilities are limited. In contrast, non-parametric models directly model the relationship between the 3D imaging ray and its resulting pixel in the image, encoding valuable priors in the learned semantic features to reason the camera parameters. Recent works [202, 203, 204] also incorporate the perspective of modeling pixel-wise information for camera calibration, making minimal assumptions on the camera model and showing more interpretable and in line with how humans perceive.
(2) Decouple different stages in an end-to-end calibration learning model. Most learning-based camera calibration methods include a feature extraction stage and an objective estimation stage. However, how the networks learn the features related to calibration is ambiguous. Therefore, decoupling the learning process by different traditional calibration stages can guide the way of feature extraction. It would be meaningful to extend the idea in extrinsic calibration [115, 167, 174] to more general calibration problems.
(3) Transfer the measurement space from the parameter error to the geometric difference. When it comes to jointly calibrating various camera parameters, the training process will suffer from an imbalance loss optimization problem. The main reason is different camera parameters correspond to different sample distributions. The simple normalization strategy cannot unify their error spaces. Therefore, we can formulate a straightforward measurement space in terms of the geometric properties of different camera parameters.
(4) Despite the strides achieved by recent NeRF methods, training NeRF without precise camera parameters remains a challenge, especially in scenarios with sparse views, pronounced movements, low-texture regions, and suboptimal initial values. Contemporary NeRF-based methodologies do optimize camera parameters, yielding notable results. However, they exhibit significant computational demands and lack the generalization seen in current deep-learning calibration techniques. We contend that in the present NeRF-based methods, camera parameters often play a secondary role. Thus, crafting effective calibration algorithms that capitalize on NeRF remains an arduous but promising endeavor.
@@ 4.1.2 Reconstruction-based Solution
Inspired by the conditional image-to-image translation and dense visual perception, the reconstruction-based solution starts to evolve from the conventional regression-based paradigm. DR-GAN [34] is the first reconstruction-based solution for calibrating the radial distortion, which directly models the pixel-wise mapping between the distorted image and rectified image. It achieved the camera parameter-free training and one-stage rectification. Thanks to the liberation of the assumption of camera models, the reconstruction-based solution showed the potential to calibrate various types of cameras in one learning network. For example, DDM [35] unified different camera models into a domain by presenting the distortion distribution map, which explicitly describes the distortion level of each pixel in a distorted image. Then, the network learned to reconstruct the rectified image using this geometric prior map. To make the mapping function interpretable, the subsequent works [96, 37, 46, 121, 145, 47, 50, 143] developed the displacement filed between the distorted image and rectified image. Such a manner is able to eliminate the generated artifacts in the pixel-level reconstruction. In particular, FE-GAN [121] integrated the geometry prior like Shi et al. [81] and PSE-GAN [119] into their reconstruction-based solution and presented a self-supervised strategy to learn the distortion flow for wide-angle camera calibration in Figure 5. Most reconstruction-based solutions exploit a U-Net-like architecture to learn pixel-level mapping. However, the distortion feature can be transferred from encoder to decoder by the skip-connection operation, leading to a blurring appearance and incomplete correction in reconstruction results. To address this issue, Li et al. [117] abandoned the skip-connection in their rectification network. To keep the feature fusion and restrain the geometric difference simultaneously, PCN [145] designed a correction layer in skip-connection and applied the appearance flows to revise the convolved features in different encoder layers. Having noticed that the previous sampling strategy of the convolution kernel neglected the radial symmetry of distortion, PolarRecNet [143] transformed the distorted image from the Cartesian coordinates domain into the polar coordinates domain.
@@ 4.2.1 Single-frame-based Solution
URS-CNN [27] is the first learning work for calibrating the rolling shutter camera. In this work, a neural network with long kernel characteristics was used to understand how the scene structure and row-wise camera motion interact. To specifically address the nature of the RS effect produced by the row-wise exposure, the row-kernel and column-kernel convolutions were leveraged to extract attributes along horizontal and vertical axes. RSC-Net [100] improved URS-CNN [27] from 2 degrees of freedom (DoF) to 6-DoF and presents a structure-and-motion-aware RS correction model, where the camera scanline velocity and depth were estimated. Compared to URS-CNN [27], RSC-Net [100] further reasoned about the concealed motion between the scanlines as well as the scene structure as shown in Figure 6. To bridge the spatiotemporal connection between RS and GS, EvUnroll [42] exploited the neuromorphic events to correct the RS effect. Event cameras can overcome a number of drawbacks of conventional frame-based activities for dynamic situations with quick motion due to their high temporal resolution property with microsecond-level sensitivity.
@@ 4.2.2 Multi-frame-based Solution
Most multi-frame-based solutions are based on the reconstruction paradigm, they mainly devote to contributing how to represent the dense displacement field between RS and global GS images and accurately warp the RS domain to the GS domain. For the first time, DeepUnrollNet [49] proposed an end-to-end network for two consecutive rolling shutter images using a differentiable forward warping module. In this method, a motion estimation network is used to estimate the dense displacement field from a rolling shutter image to its matching global shutter image. The second contribution of DeepUnrollNet [49] is to construct two novel datasets: the Fastec-RS dataset and the Carla-RS dataset. Furthermore, JCD [153] jointly considered the rolling shutter correction and deblurring (RSCD) techniques, which largely exist in the medium and long exposure cases of rolling shutter cameras. It applied bi-directional warping streams to compensate for the displacement while keeping the non-warped deblurring stream to restore details. The authors also contributed a real-world dataset using a well-designed beam-splitter acquisition system, BS-RSCD, which includes both ego-motion and object motion in dynamic scenes. SUNet [159] extended DeepUnrollNet [49] from the middle time of the second frame () into the intermediate time of two frames (). By using PWC-Net [205], SUNet [159] estimated the symmetric undistortion fields and reconstructed the potential GS frames by a time-centered GS image decoder network. To effectively reduce the misalignment between the contexts warped from two consecutive RS images, the context-aware undistortion flow estimator and the symmetric consistency enforcement were designed. To achieve a higher frame rate, Fan et al. [158] generated a GS video from two consecutive RS images based on the scanline-dependent nature of the RS camera. In particular, they first analyzed the inherent connection between bidirectional RS undistortion flow and optical flow, demonstrating the RS undistortion flow map has a more pronounced scanline dependency than the isotropically smooth optical flow map. Then, they developed the bidirectional undistortion flows to describe the pixel-wise RS-aware displacement, and further devised a computation technique for the mutual conversion between different RS undistortion flows corresponding to various scanlines. To eliminate the inaccurate displacement field estimation and error-prone warping problems in previous methods, AW-RSC [166] proposed to predict multiple fields and adaptively warped the learned RS features into global shutter counterparts. Using a coarse-to-fine approach, these warped features were combined and generated to precise global shutter frames as shown in Figure 7. Compared to previous works [49, 153, 159, 158], the warping operation consisting of adaptive multi-head attention and a convolutional block in AW-RSC [166] is learnable and effective. In addition, AW-RSC [166] contributed a real-world rolling shutter correction dataset: BS-RSC, where the RS videos with corresponding GS ground truth are captured simultaneously with a beam-splitter-based acquisition system.
@@ 4.3.1 Technique Summary
The deep learning works on wide-angle camera and roll shutter calibration share a similar technique pipeline. Along this research trend, most early literature begins with the regression-based solution [26, 40, 27]. The subsequent works innovated the traditional calibration with a reconstruction perspective [34, 35, 121, 49], which directly learns the displacement field to rectify the uncalibrated input. For higher accuracy of calibration, a more intuitive displacement field, and more effective warping strategy have been developed [145, 166, 153, 158]. To fit the distribution of different distortions, some works designed different shapes of the convolutional kernel [27] or transformed the convolved coordinates [143].
Existing works devoted themselves to designing more powerful networks and introducing more diverse features to facilitate calibration performance. Increasingly more methods focused on the geometry priors of the distortion [121, 119, 81]. These priors can be directly weighted into the convolutional layers or used to supervise network training, promoting the learning model to converge faster.
@@ 4.3.2 Future Effort
(1) The development of wide-angle camera calibration and roll shutter camera calibration can promote each other. For instance, the well-studied multi-frame-based solution in roll shutter calibration is able to inspire wide-angle calibration. The same object located at different sequences could provide useful priors regarding to radial distortion. Additionally, the elaborate studies of the displacement field and warping layer [166, 153, 158] have the potential to motivate the development of wide-angle camera calibration and other fields. Furthermore, the investigation of geometric priors in wide-angle calibration could also improve the interpretability of the network in roll shutter calibration.
(2) Most methods synthesize their training dataset based on random samples from all camera parameters. However, for the images captured by real lenses, the distribution of camera parameters probably locates at a potential manifold [88]. Learning on a label-redundant calibration dataset makes the training process inefficient. Thus, exploring a practical sampling strategy for the synthesized dataset could be a meaningful task in the future direction.
(3) To overcome the ill-posed problem of single-frame calibration, introducing other high-precision sensors can compensate for the current calibration performance, such as event cameras [42]. With the rapid development of vision sensors, joint calibration using multiple sensors is valuable. Consequently, more cross-modal and multi-modal fusion techniques will be investigated along this research way.
@@ 5 Cross-View Model
The existing deep calibration methods can estimate the specific camera parameters from a single camera. In fact, there can be more complicated parameter representations in multi-camera circumstances. For example, in the multi-view model, the fundamental matrix and essential matrix describe the epipolar geometry and they are intricately tangled with intrinsics and extrinsics. The homography depicts the pixel-level correspondences between different views. In addition to intrinsics and extrinsics, it is also intertwined with depth. Among these complex parameter representations, homography is the most widely leveraged in practical applications and its related learning-based methods are the most investigated. To this end, we mainly focus on the review of deep homography estimation solutions for the cross-view model and they can be divided into three categories: direct, cascaded, and iterative solution.
@@ 5.1 Direct Solution
Then the 4-pt parameterization can be solved as a homography matrix using normalized DLT[206]. However, DHN is limited to synthetic datasets where the ground truth can be generated for free or requires costly labeling of real-world datasets. Subsequently, the first unsupervised solution named UDHN[52] is proposed to address this problem. As shown in Fig. 8(c), it used the same network architecture as DHN and defined an unsupervised loss function by minimizing the average photometric error motivated by traditional methods[207]:
@@ 5.2 Cascaded Solution
Direct solutions explore various homography parameterizations with simple network structures, while the cascaded ones focus on complex designs of network architectures.
In HierarchicalNet[72], Nowruzi et al. hold that the warped images can be regarded as the input of another network. Therefore they stacked the networks sequentially to reduce the error bounds of the estimate. Based on HierarchicalNet, SRHEN [136] introduced the cost volume[205] to the cascaded network, measuring the feature correlation by cosine distance and formulating it as a volume. The stacked networks and cost volume increase the performance, but they cannot handle the dynamic scenes. MHN [113] developed a multi-scale neural network and proposed to learn homography estimation and dynamic content detection simultaneously. Moreover, to tackle the cross-resolution problem, LocalTrans [147] formulated it as a multimodal problem and proposed a local transformer network embedded within a multiscale structure to explicitly learn correspondences between the multimodal inputs. These inputs include images with different resolutions, and LocalTrans achieved superior performance on cross-resolution cases with a resolution gap of up to 10x. All the solutions mentioned above leverage image pyramids to progressively enhance the ability to address large displacements. However, every image pair at each level requires a unique feature extraction network, resulting in the redundancy of feature maps. To alleviate this problem, some researchers[210, 149, 44, 55] replaced image pyramids with feature pyramids. Specifically, they warped the feature maps directly instead of images to avoid excessive feature extraction networks. To address the low-overlap homography estimation problem in real-world images[149], Nie et al.[149] modified the unsupervised constraint (Eq. 3) to adapt to low-overlap scenes:
where is an all-one matrix with the same size as or . It solved the low-overlap problem by taking the original images as network input and ablating the corresponding pixels of to the invalid pixels of warped . To solve the non-planar homography estimation problem, DAMG-Homo[44] proposed backward multi-gird deformation with contextual correlation to align parallax images. Compared with traditional cost volume, the proposed contextual correlation helped to reach better accuracy with lower computational complexity. Another way to address the non-planar problem is to focus on the dominant plane. In HomoGAN [55], an unsupervised GAN is proposed to impose a coplanarity constraint on the predicted homography, as shown in Figure 9. To implement this approach, a generator is used to predict masks of aligned regions, while a discriminator is used to determine whether two masked feature maps were produced by a single homography.
@@ 5.3 Iterative Solution
where and are the feature maps of the template and input images. Then, they enforced the network to run a single iteration with a hinge loss, while the network runs multiple iterations until the stopping condition is met in the testing stage. Besides, CLKN stacked three similar LK networks to further boost the performance by treating the output of the last LK network as the initial warp parameters of the next LK network. From Eq. 7, the IC-LK algorithm heavily relied on feature maps, which tend to fail in multimodal images. Instead, DLKFM [146] constructed a single-channel feature map by using the eigenvalues of the local covariance matrix on the output tensor. To learn DLKFM, it designed two special constraint terms to align multimodal feature maps and contribute to convergence.
However, LK-based algorithms can fail if the Jacobian matrix is rank-deficient [212]. Additionally, the IC-LK iterator is untrainable, which means this drawback is theoretically unavoidable. To address this issue, a completely trainable iterative homography network (IHN) [165] was proposed. Inspired by RAFT [213], IHN updates the cost volume to refine the estimated homography using the same estimator repeatedly every iteration. Furthermore, IHN can handle dynamic scenes by producing an inlier mask in the estimator without requiring extra supervision.

### v3

@@ 3.4.2 Future Effort
(1) Explore more model priors. Most learning-based methods study the parametric camera models but their generalization abilities are limited. In contrast, non-parametric models directly model the relationship between the 3D imaging ray and its resulting pixel in the image, encoding valuable priors in the learned semantic features to reason the camera parameters. A non-parametric model based on implicit neural representations, as explored in [363], has shown remarkable advantages in accurately capturing the camera within a higher-dimensional space. This approach holds potential for future extensions, such as modeling physical factors like aberration. Recent works [175, 172, 165] incorporate the perspective of modeling pixel-wise information for camera calibration, making minimal assumptions on the camera model and showing more interpretable and in line with how humans perceive.
(2) Decouple different stages in an end-to-end calibration learning model. Most learning-based methods include a feature extraction stage and an objective estimation stage. However, how the networks learn the features related to calibration is ambiguous. Therefore, decoupling the learning process by different traditional calibration stages can guide the way of feature extraction. It would be meaningful to extend the idea from the extrinsic calibration [112, 146, 153] to more general calibration problems.
(3) Transfer the measurement space from the parameter error to the geometric difference. Jointly calibrating multiple camera parameters poses an optimization imbalance due to their varying sample distributions. Simple normalization fails to unify their error spaces. A potential solution is to establish a direct measurement space based on the geometric properties of different camera parameters.
(4) Training NeRF without precise camera parameters remains challenging, particularly in scenarios with sparse views, significant motion, low-texture regions, and suboptimal initial values. While modern NeRF-based methods optimize camera parameters with notable results, they are computationally demanding and lack the generalization of deep-learning calibration techniques. We argue that in current NeRF-based approaches, camera parameters often play a secondary role. Therefore, developing effective calibration algorithms that leverage NeRF remains a difficult yet promising pursuit.
@@ 4.1.2 Reconstruction-based Solution
Inspired by the image-to-image translation and dense visual perception, the reconstruction-based solution starts to evolve from the conventional regression-based paradigm. DR-GAN [35] is the first reconstruction-based solution for calibrating the radial distortion, which directly models the pixel-wise mapping between the distorted image and the rectified image. It achieved the camera parameter-free training and one-stage rectification. Thanks to the liberation of the assumption of camera models, the reconstruction-based solution showed the potential to calibrate various types of cameras in one learning network. For example, DDM [36] unified different camera models into a domain by presenting the distortion distribution map, which explicitly describes the distortion level of each pixel in an image. Then, the network learned to reconstruct the rectified image using this geometric prior map. To make the mapping function interpretable, the subsequent works [97, 38, 52, 120, 136, 53, 56, 134, 181, 224] developed the displacement field between the distorted image and rectified image. Such a manner eliminates the generated artifacts in the pixel-level reconstruction. In particular, FE-GAN [120] integrated the geometry prior like Shi et al. [85] and PSE-GAN [118] into their reconstruction-based solution and presented a self-supervised strategy to learn the distortion flow for wide-angle camera calibration in Figure 5. Most reconstruction-based solutions exploit a U-Net-like architecture to learn pixel-level mapping. However, the distortion feature can be transferred from encoder to decoder by the skip-connection operation, leading to a blurring appearance and incomplete correction in reconstruction results. To address this issue, Li et al. [116] abandoned the skip-connection in their rectification network. To keep the feature fusion and restrain the geometric difference simultaneously, PCN [136] designed a correction layer in skip-connection and applied the appearance flows to revise the features in different encoder layers. Having noticed that the previous sampling strategy of the convolution kernel neglected the radial symmetry of distortion, some works [134, 364] transformed the image from the Cartesian coordinates domain into the polar coordinates domain.
@@ 4.2.1 Single-frame-based Solution
URS-CNN [28] is the first learning work for calibrating the rolling shutter camera. In this work, a neural network with long kernel characteristics was used to understand how the scene structure and row-wise camera motion interact. To specifically address the nature of the RS effect produced by the row-wise exposure, the row-kernel and column-kernel convolutions were leveraged to extract attributes along horizontal and vertical axes. RSC-Net [99] improved URS-CNN [28] from 2 degrees of freedom (DoF) to 6-DoF and presents a structure-and-motion-aware model, where the camera scanline velocity and depth were estimated. Compared to URS-CNN [28], RSC-Net [99] reasoned about the concealed motion between the scanlines as well as the scene structure as shown in Figure 6. To bridge the spatiotemporal connection between RS and GS, some methods [47, 228, 187, 229] exploited the neuromorphic events to correct the RS effect. Event cameras can overcome many drawbacks of conventional frame-based activities for dynamic situations with quick motion due to their high temporal resolution property with microsecond-level sensitivity.
@@ 4.2.2 Multi-frame-based Solution
Most multi-frame-based solutions are based on the reconstruction paradigm, they mainly devote to contributing how to represent the dense displacement field between RS and global GS images and accurately warp the RS domain to the GS domain. For the first time, DeepUnrollNet [55] proposed an end-to-end network for two consecutive rolling shutter images using a differentiable forward warping module. In this method, a motion estimation network is used to estimate the dense displacement field from a rolling shutter image to its matching global shutter image. The second contribution of DeepUnrollNet [55] is to construct two novel datasets: the Fastec-RS dataset and the Carla-RS dataset. Furthermore, JCD [137] jointly considered the rolling shutter correction and deblurring (RSCD) techniques, which largely exist in the medium and long exposure cases of rolling shutter cameras. It applied bi-directional warping streams to compensate for the displacement while keeping the non-warped deblurring stream to restore details. The authors also contributed a real-world dataset using a well-designed beam-splitter acquisition system, BS-RSCD, which includes both ego-motion and object motion in dynamic scenes. SUNet [140] extended DeepUnrollNet [55] from the middle time of the second frame () into the intermediate time of two frames (). By using PWC-Net [365], SUNet [140] estimated the symmetric undistortion fields and reconstructed the potential GS frames by a time-centered GS image decoder network. To effectively reduce the misalignment between the contexts warped from two consecutive RS images, the context-aware undistortion flow estimator and the symmetric consistency enforcement were designed. To achieve a higher frame rate, Fan et al. [139] generated a GS video from two consecutive RS images based on the scanline-dependent nature of the RS camera. In particular, they first analyzed the inherent connection between bidirectional RS undistortion flow and optical flow, demonstrating the RS undistortion flow map has a more pronounced scanline dependency than the isotropically smooth optical flow map. Then, they developed the bidirectional undistortion flows to describe the pixel-wise RS-aware displacement, and further devised a computation technique for the mutual conversion between different RS undistortion flows corresponding to various scanlines. To eliminate the inaccurate displacement field estimation and error-prone warping problems in previous methods, AW-RSC [158] proposed to predict multiple fields and adaptively warped the learned RS features into global shutter counterparts. Using a coarse-to-fine approach, these warped features were combined and generated to precise global shutter frames as shown in Figure 7. Compared to previous works [55, 137, 140, 139], the warping operation consisting of adaptive multi-head attention and a convolutional block in AW-RSC [158] is learnable and effective. The following works further incorporate more tractable designs for accurate RS corrections, such as directly estimating the Distortion Flow from consecutive RS frames [230], jointly performing the GS appearance refinement and correction motion estimation [193], jointly learning rolling shutter correction and super-resolution [192], and introducing a self-supervised learning framework with bidirectional distortion warping modules [188], etc.
@@ 4.3.1 Technique Summary
The deep learning works on wide-angle camera and roll shutter calibration share a similar technique pipeline. Along this research trend, most early literature begins with the regression-based solution [27, 45, 28]. The subsequent works innovated the traditional calibration with a reconstruction perspective [35, 36, 120, 55], which directly learns the displacement field to rectify the uncalibrated input. For higher accuracy of calibration, a more intuitive displacement field, and more effective warping strategy have been developed [136, 158, 137, 139, 195]. To fit the distribution of different distortions, some works designed different shapes of the convolutional kernel [28], transformed the convolved coordinates [134], and equipped with the distortion-aware modules [48, 186, 364].
Existing works devoted themselves to designing more powerful networks and introducing more diverse features to facilitate calibration performance. Recent methods focused on the geometry priors of the distortion [120, 118, 85, 219, 221, 175]. These priors can be directly weighted into the convolutional layers or used to supervise network training, promoting the learning model to converge faster.
@@ 4.3.2 Future Effort
(1) The development of wide-angle camera calibration and roll shutter camera calibration could promote each other. For instance, the well-studied multi-frame-based solution in roll shutter calibration is able to inspire wide-angle calibration. The same object located at different sequences could provide useful priors regarding to radial distortion. Additionally, the elaborate studies of the displacement field and warping layer [158, 137, 139] have the potential to motivate the development of wide-angle camera calibration and other fields. Furthermore, the investigation of geometric priors in wide-angle calibration could also improve the interpretability of the network in roll shutter calibration.
(2) Most methods synthesize their training dataset based on random samples from all camera parameters. However, for the images captured by real lenses, the distribution of camera parameters is probably located at a potential manifold [93]. Learning on a label-redundant calibration dataset makes the training process inefficient. Thus, exploring a practical sampling strategy for the synthesized dataset could be a meaningful task in the future direction.
(3) To overcome the ill-posed problem of single-frame calibration, introducing other high-precision sensors can compensate for the current calibration performance, such as event cameras [47, 190, 191, 187, 229]. With the rapid development of vision sensors, joint calibration using multiple sensors is highly valuable. Consequently, more cross-modal and multi-modal fusion techniques will be investigated along with this research path.
@@ 5 Cross-View Model
Existing methods can estimate specific camera parameters for a single camera. However, in multi-camera scenarios, parameter representations can be more complex. For example, in the multi-view model, the fundamental matrix and essential matrix describe the epipolar geometry and they are intricately tangled with intrinsics and extrinsics [261, 284, 286]. The homography depicts the pixel-level correspondences between different views. In addition to intrinsics and extrinsics, it is also intertwined with depth. Among these complex parameter representations, homography is the most widely leveraged and its related learning-based methods are the most investigated. To this end, we mainly focus on the review of homography estimation solutions for the cross-view model. They can be divided into three categories: direct, cascaded, and iterative solutions.
@@ 5.1 Direct Solution
The 4-pt parameterization can be solved as a homography matrix using normalized DLT[366]. However, DHN is limited to synthetic datasets where the ground truth can be generated for free or requires costly labeling of real-world datasets. Subsequently, the unsupervised solution named UDHN[58] is proposed to address this problem. As shown in Fig. 8(c), it used the same network architecture as DHN and defined an unsupervised loss function by minimizing the average photometric error motivated by traditional methods[367]:
@@ 5.2 Cascaded Solution
Direct solutions explore various homography parameterizations with simple network structures, while the cascaded ones focus on complex designs of network architectures. HierarchicalNet[232] held that the warped images can be regarded as the input of another network. Therefore it stacked the networks sequentially to reduce the error bounds of the estimate. Based on HierarchicalNet, SRHEN [241] introduced the cost volume[365] to the cascaded network, measuring the feature correlation by cosine distance and formulating it as a volume. The stacked networks and cost volume increase the performance, but they cannot handle the dynamic scenes. MHN [240] developed a multi-scale neural network and proposed to learn homography estimation and dynamic content detection simultaneously. Moreover, to tackle the cross-resolution problem, LocalTrans [249] formulated it as a multimodal problem and proposed a local transformer network embedded within a multiscale structure to explicitly learn correspondences between the multimodal inputs. These inputs include images with different resolutions, and LocalTrans achieved superior performance on cross-resolution cases with a resolution gap of up to 10. All the solutions mentioned above leverage image pyramids to progressively enhance the ability to address large displacements. However, every image pair at each level requires a unique feature extraction network, resulting in the redundancy of feature maps. To alleviate this problem, some researchers[370, 251, 49, 61] replaced image pyramids with feature pyramids. Specifically, they warped the feature maps directly instead of images to avoid excessive feature extraction networks. To address the low-overlap homography estimation problem in real-world images[251], Nie et al.[251] modified the unsupervised constraint (Eq. 3) to adapt to low-overlap scenes:
where is an all-one matrix with the same size as or . It solved the low-overlap problem by taking the original images as network input and ablating the corresponding pixels of to the invalid pixels of warped . To solve the non-planar homography estimation problem, DAMG-Homo[49] proposed backward multi-gird deformation with contextual correlation to align parallax images. Compared with traditional cost volume, the proposed contextual correlation helped to reach better accuracy with lower computational complexity. Another way to address the non-planar problem is to focus on the dominant plane. In HomoGAN [61], an unsupervised GAN is proposed to impose a coplanarity constraint on the predicted homography, as shown in Figure 9. Specifically, a generator is used to predict masks of aligned regions, while a discriminator is used to determine whether two masked feature maps were produced by a single homography. To eliminate the lack of adequate labeled training data, recent methods explored generating realistic and accurate labels using diffusion models [289] and dominant plane detection network [266], etc.
@@ 5.3 Iterative Solution
where and are the feature maps of the template and input images. Then, they enforced the network to run a single iteration with a hinge loss, while the network runs multiple iterations until the stopping condition is met in the testing stage. Besides, CLKN stacked three similar LK networks to boost the performance by treating the output of the last LK network as the initial warp parameters of the next LK network. From Eq. 7, the IC-LK algorithm heavily relied on feature maps, which tend to fail in multimodal images. Instead, DLKFM [248] constructed a single-channel feature map using the eigenvalues of the local covariance matrix on the output tensor. It designed two special constraint terms to align multimodal feature maps and contribute to convergence.
However, LK-based algorithms can fail if the Jacobian matrix is rank-deficient [372]. Additionally, the IC-LK iterator is untrainable, which means this drawback is theoretically unavoidable. To address this issue, a completely trainable iterative homography network (IHN) [255] was proposed. Inspired by RAFT [373], IHN updates the cost volume to refine the estimated homography using the same estimator repeatedly every iteration. Furthermore, IHN can handle dynamic scenes by producing an inlier mask in the estimator without requiring extra supervision. The following works further explore more efficient and holistic frameworks for the cross-modal homography estimation [309, 310, 305, 298].